OpenStudy (highschoolmom2010):

write the geometric mean 5,1.25

OpenStudy (rsadhvika):

multiply them, take square root

OpenStudy (highschoolmom2010):

6.25

OpenStudy (highschoolmom2010):

\[\sqrt{6.25}\] cant be reduced

OpenStudy (rsadhvika):

it can be reduced

OpenStudy (rsadhvika):

notice that 625 = 25^2

OpenStudy (rsadhvika):

\(\large \sqrt{6.25}\)

OpenStudy (highschoolmom2010):